OpenStudy (anonymous):

Find the x- and y- intercepts. f(x)=9x^+24x+16 x= y=

OpenStudy (anonymous):

I think you wrote the equation wrong

OpenStudy (anonymous):

no it is written correctly

OpenStudy (anonymous):

you have x to the power of [blank]

OpenStudy (anonymous):

F(x)=9x²+24x+16

OpenStudy (anonymous):

this should help

OpenStudy (anonymous):

The y intercept is the value of y when x is 0, so replace x with 0 and solve for y:\[y = 9(0)^{2}+24(0)+16\]\[y=16\]The x intercept is the x value when y is 0, so replace y with 0 and solve for x:\[0 = 9x^{2}+24x+16\]\[0=(3x+4)^{2}\]\[0=3x+4\]\[-\frac{4}{3}=x\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so if written as an ordered pair for y it would be (0,16) and for x it would be (0,-4/3) correct.

OpenStudy (anonymous):

No, for the x-intercept (the 2nd ordered pair), it would be (-4/3, 0)

OpenStudy (anonymous):

ok, I just wanted to make sure. Could not remember and that is why I asked

OpenStudy (anonymous):

It looks like this: http://www.wolframalpha.com/input/?i=y+%3D+9x%5E2%2B24x%2B16&cdf=1

OpenStudy (anonymous):

thank you so much :)